### v2

# Does Quantum Mechanics apply to Macroscopic Objects? How to define ”Macro”?

Abstract
According to Quantum Mechanics a narrow wave-packet of the center of mass of any macroscopic object should smear out after some time. The problem is usually waved out by assuming that due to their heavy masses this occurs over astronomical times. Without a clear definition of macroscopic objects this remains ambiguous. On the other hand, Quantum Mechanics allows and energetically even prefers largely smeared out c.m. states that have never been seen. Why is this the real state of the world we know? Does it suggest a micro-macro threshold with different theoretical descriptions? Does micro-miniaturization offer hope for answering these questions?
Institut für Theoretische Physik, Goethe-Universität, Frankfurt am Main
banyai@itp.uni-frankfurt.de corresponding author
ORCID (https://www.orcid.org/$0$-0002-0698-4202)
Keywords: Quantum Mechanics; Macroscopic Objects; Center of Mass; Wave Packets; Condensed Matter; Micro-Miniaturization, Quantum Domain
1 Introduction
One of the strange features of Quantum Mechanics is that for its formulation, one needs the classical physics that one expects to emerge as its macroscopic limit. All experiences with quantum objects have to be analysed through classical ”glasses”. On the other hand, classical theory describes well the behaviour of macroscopic objects around us. But what is the actual definition of a macroscopic object? Where is the threshold (if any) between the microscopic and the macroscopic worlds?
One should bear in mind that Quantum Mechanics is just a theoretical frame to describe the properties of a system that may be treated as being isolated and having a finite number of degrees of freedom. Its basic elements are the concept of the wave function (quantum state), the Schrödinger equation for its evolution, the observables and the probabilistic interpretation of the measurement. The system itself must be defined first from the experimental point of view and this is based on our general vague conception and long experience about the surrounding world. Then one associates a certain Hamiltonian to it and verifiable mathematical predictions of Quantum Mechanics emerge. Strictly speaking, there are no isolated physical systems and the number of degrees of freedom of a system depends on our modelling it.
Quantum Mechanics says nothing about its object or the structure of matter. It is an exceptionally successful theory, nevertheless the relationship between the macroscopic classical and the microscopic quantum mechanical description of matter is still subject of worry.
In the first half of the twentieth century the big successes of quantum theory were obtained mainly in the field of (”elementary”) microscopic particles: electrons, atoms, nuclei. Thereafter, the most spectacular evolution occurred in understanding the properties of condensed matter based on many-body quantum mechanics. This step revolutionized our technology.
Now we witness another exceptional development in producing ever smaller pieces of solid matter (semiconductor chips, micro-structures). We are approaching the microscopic world from above. Therefore, the answer to the question about the upper end of the ”quantum-domain”, might be very close.
I want here to underline some simple but often ignored aspects in this context. The theory of condensed matter is conceived as the quantum mechanical description of bound states of an enormous number of Coulomb interacting electrons and ions. This picture is very successful. In the absence of an external potential the center of mass (c.m.) motion is fully separable from the relative (internal) one [1] and its wave function satisfies the free Schrödinger equation of a particle with the total mass $M$ of the system. We apply Quantum Mechanics however, only to the relative motion, while the center of mass motion is considered irrelevant and left to the classical description.
In what follows we shall discuss only the center of mass motion. In the real world the center of mass states of the macroscopic objects surrounding us are from the quantum mechanical point of view sharply defined wave packets. According to the Schrödinger equation these should spread out in time. A somewhat appeasing argument is given in the frame of Ehrenfest’s theorem in almost all handbooks about their stability for macroscopic (heavy) masses. However, no satisfactory definition of the term ”macroscopic” was given until now.
In Section 2 we recapitulate the standard quantum mechanical description of an ideal wave packet (here the free motion of the c.m), while in Section 3 we comment on the not understood absence of broad wave packets of macroscopic objects in the real word.
2 Evolution of a Wave-Packet of the c.m.
Let us consider the solution of the free 1D Schrödinger equation for an object of mass $M$ in the form of a Gaussian wave packet (having the ideal uncertainty $\Delta x\Delta p=\hbar$) travelling with a constant average velocity. It may represent the free motion of the c.m. of a macroscopic object of total mass $M$. Using the Galilean invariance of the free Schrödinger equation [2] one may always go to its own frame (in our case the c.m. frame of reference) defined by the vanishing of the average velocity and at some initial time $t=0$ one has the wave function
$\psi(x,0)=\frac{1}{{(2\pi)}^{\frac{1}{4}}\sqrt{d}}e^{-\frac{x^{2}}{4d^{2}}}\qquad.$
Then after a lapse of time $t$ it decays
$\psi(x,t)=\frac{1}{(2\pi)^{\frac{1}{4}}\sqrt{d+\frac{\imath\hbar t}{2dM}}}e^{-\frac{x^{2}}{4(d^{2}+\frac{\imath\hbar t}{2M})}}\qquad.$
This means that the average quadratic width grows as
$\langle x^{2}\rangle_{t}=d^{2}+(\frac{\hbar t}{2Md})^{2}\qquad.$
Of course, the discussion may be extended to the motion in the presence of a slowly varying external potential on the scale of the macroscopic object as in Ehrenfest’s Theorem. All this is well-known.
One may see, that for $M\to\infty$ the width does not change at all. However, the decisive parameter is not the mass $M$ alone but the product $Md$. For any finite mass $M$ at $d\to 0$ the wave packet decays instantly.
[…]
No doubt, the c.m. of a hydrogen atom obeys Quantum Mechanics, but does it obey also the c.m of this small crystal? The answer seems to be negative. Then how do we define a ”macroscopic” object?
3 What about the ”initial” state?
The next provoking question is: Why all macroscopic objects we know have very precise center of mass position? Ultimately, why are there no plane waves of macroscopic objects, although energetically more favourable. In our example of the Gaussian packet the average kinetic energy is
$\frac{\hbar^{2}}{2M}\langle k^{2}\rangle=\frac{\hbar^{2}}{8Md^{2}}\qquad.$
For non-ideal wave-packets $\Delta x\Delta p>\hbar$ the average kinetic energy is even greater. Therefore the more smeared is the wave packet, the smaller its conserved average energy is. Although energetically preferred, no smeared out macroscopic objects have ever been seen. It looks like a ”super-selection” rule for macroscopic objects. ”God allowed them only in this state of precise c.m.?” On the other hand, a world with smeared out mass centers is hard to imagine. It won’t function. The macroscopic world, as we know it, is the only possible world!
There is another striking old problem about the application of quantum-mechanical concepts to macroscopic objects. There are no known superpositions of states of macroscopic systems. It was illustrated by the famous ”Schrödinger’s cat”.
One arrives at the conclusion that we have two different worlds: microscopic and macroscopic ones , with different theoretical descriptions.
In the above sense, I would like to encourage experiments about the imprecision of the c.m of ultra small pieces of crystals. On our current way toward micro-miniaturization we might get an answer to the question about the threshold between micro and macro. It is very plausible that the macroscopic physics does not emerge from the microscopic theory. (In a mathematical sense it is surely not the $\hbar\to 0$ limit.)
This should solve also the ”bootstrap aspect” of Quantum Mechanics i.e. the existence of macroscopic laws as prerequisite for its formulation and application. In other words: Quantum Mechanics has its limited field of validity as any other known physical theory.
4 Comments
I am aware that just touching this point might look like a heresy. In our human-centred world-image motivated by the exceptional scientific progresses of the last centuries we cheer to construct a unified picture of the world without loopholes and crevices. One looks even for ”God’s equations” forgetting that mathematics has to be connected to experiment by some interpretation and this is the most difficult one.
However, all theories of physics of the past were just fragments of knowledge (models) with loose connections between them. Thermodynamics was built up much before statistical mechanics and their link is still rather vague. In the modern functional integral formalism of quantum field theory of elementary particles there is no place for wave functions, Hamiltonians and observables while its mathematical structure is highly questionable. Nevertheless, it succeeded to predict the fauna of elementary particles. However, its connection to the ordinary quantum mechanics of systems with finite degrees of freedom remains a question of belief (see also [3]). The unification of gravitation and quantum theory looks merely a dream.
Theoretical physics never had a strict vertical structure. It was never derived from the ”underlying” theories of matter. I am afraid, that we are misled by our need for harmony and exaggerated confidence in our brains.
[…]